Analyze exercise 12-23 on page 619. Calculate (a) the partnership's gain of loss on realization and (b) how much cash each partner will receive after debts are paid and the firm is liquidated. Show your work.

12-23

After closing the accounts on July 1, prior to liquidating the partnership, the capital ac-
count balances of Gold, Porter, and Sims are $55,000, $45,000, and $20,000, respectively. Cash, noncash assets, and liabilities total $56,000, $96,000, and $32,000, respectively. Between July 1 and July 29, the noncash assets are sold for $90,000, the liabilities are paid, and the remaining cash is distributed to the partners. The partners share net in- come and loss in the ratio of 3:2:1. Prepare a statement of partnership liquidation for the period July 1–29.
Answer #1

Answer

  • Statement of Partnership liquidation, for period Jul 1 – 29

Cash

Non Cash Asset

Liabilities

Gold, Capital

Porter, Capital

Sims, Capital

Beginning balances

$                56,000

$               96,000

$               32,000

$               55,000

$              45,000

$                     20,000

     Non Cash Asset sold

$                90,000

$             (96,000)

$                        -  

$               (3,000)

$              (2,000)

$                     (1,000)

Balance after non cash assets are sold and losses distributed

$              146,000

$                        -  

$               32,000

$              52,000

$              43,000

$                     19,000

     Liabilities paid off

$              (32,000)

$                        -  

$            (32,000)

$                        -  

$                       -   

$                               -  

Balances after liabilities were paid off

$              114,000

$                        -  

$                        -  

$               52,000

$              43,000

$                     19,000

     Final balance cash distributed

$            (114,000)

$                        -  

$                        -  

$             (52,000)

$            (43,000)

$                   (19,000)

Ending balance after liquidation

$                          -  

$                        -  

$                        -  

$                        -  

$                       -  

$                               -  

  • Working for Loss on non cash asset sold:

--Value = $ 96000, Sold for $ 90000. Loss = 96000 – 90000 = $ 6000
--Loss distributed to:
Gold = 6000 x 3/6 = $ 3000
Porter = 6000 x 2/6 = $ 2000
Sims = 6000 x 1/6 = $ 1000